OpenStudy (anonymous):

Verify the identity. Justify each step. tan 0 + cot 0 = 1/sin 0 cos 0

ganeshie8 (ganeshie8):

those 0 s are \(\theta\) ?

OpenStudy (anonymous):

yea i just didnt know how to type them in sorry.

OpenStudy (raden):

this θ , for you ;)

OpenStudy (anonymous):

tan 0 = sin0/cos0 cot 0 = cos0/sin0 Cross multiplying sin0sin0+cos0cos0 / cos0sin0 = sin^2 0+cos^2 0 /cos0sin0 Using the identity sin^2 0+cos^2 0 = 1 =1/ cos0sin0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yea alot thanks